# Certain finite groups whose commuting conjugacy class graph satisfy Hansen-Vukičević conjecture

### v1

@@ Abstract.
In this paper we compute first and second Zagreb indices of commuting conjugacy class graphs of finite groups and determine several classes of finite groups such that their commuting conjugacy class graphs satisfy Hansen-Vukičević conjecture.
@@
*Corresponding author
@@ 1. Introduction
Let $\mathfrak{G}$ be the set of all graphs. A topological index is a function $T:\mathfrak{G}\to\mathbb{R}$ such that $T(\Gamma_{1})=T(\Gamma_{2})$ whenever the graphs $\Gamma_{1}$ and $\Gamma_{2}$ are isomorphic. Initially, topological indices were used to describe several chemical properties of molecules. Various topological indices have been defined since 1947 by using different parameters of graphs. Wiener index is the oldest topological index, introduced by Wiener [30], and it is a distance based index. Among the degree-based topological indices, the Zagreb indices are the first and were introduced by Gutman and Trinajstić [14] in 1972. They were used in examining the dependence of total $\pi$-electron energy on molecular structure. As noted in [22], Zagreb indices are also used in studying molecular complexity, chirality, ZE-isomerism and heterosystems etc. Later on, general mathematical properties of various topological indices are also studied by many mathematicians. A survey on mathematical properties of Zagreb indices can be found in [15].
Let $\Gamma$ be a simple undirected graph with vertex set $v(\Gamma)$ and edge set $e(\Gamma)$. The first and second Zagreb indices of $\Gamma$, denoted by $M_{1}(\Gamma)$ and $M_{2}(\Gamma)$ respectively, are defined as
$M_{1}(\Gamma)=\sum\limits_{v\in v(\Gamma)}\deg(v)^{2}\text{ and }M_{2}(\Gamma)=\sum\limits_{uv\in e(\Gamma)}\deg(u)\deg(v),$
where $\deg(v)$ is the number of edges incident on $v$ (called degree of $v$). Comparing first and second Zagreb indices, Hansen and Vukičević [16] posed the following conjecture in 2007.
@@
(Hansen-Vukičević Conjecture) For any simple finite graph $\Gamma$,
@@
It was shown in [16] that the conjecture is not true if $\Gamma=K_{1,5}\sqcup K_{3}$. However, Hansen and Vukičević [16] showed that Conjecture 1.1 holds for chemical graphs. In [28], it was shown that the conjecture holds for trees with equality in (1) when $\Gamma$ is a star graph. In [18], it was shown that the conjecture holds for connected unicyclic graphs with equality when the graph is a cycle. The case when equality holds in (1) is studied extensively in [29]. A survey on comparing Zagreb indices can be found in [19]. However, the search of graphs validating or invalidating Conjecture 1.1 is not completed yet. Recently, Das et al. [8] have obtained various finite non-abelian groups such that their commuting/non-commuting graphs satisfy Hansen-Vukičević Conjecture. Das and Nath [9] also obtained several finite groups such that their super commuting graphs satisfy Hansen-Vukičević Conjecture. In this paper, we compute first and second Zagreb indices of commuting conjugacy class graphs of finite groups and determine several classes of finite groups such that their commuting conjugacy class graphs satisfy Hansen-Vukičević conjecture. In Section 2.1, we consider commuting conjugacy class graphs of dihedral groups $D_{2m}$ presented by $\langle a,b:a^{m}=b^{2}=1,bab^{-1}=a^{-1}\rangle$ for $m\geq 3$, dicyclic groups $Q_{4m}$ presented by $\langle a,b:a^{2m}=1,b^{2}=a^{n},bab^{-1}=a^{-1}\rangle$ for $m\geq 2$, semidihedral groups $SD_{8m}$ presented by $\langle a,b:a^{4m}=b^{2}=1,bab=a^{2m-1}\rangle$ for $m\geq 2$, the group $V_{8m}$ presented by $\langle a,b:a^{2m}=b^{4}=1,ba=b^{-1}a^{-1},b^{-1}a=a^{-1}b\rangle$ for $m\geq 1$, the group $U_{(n,m)}$ presented by $\langle a,b:a^{2n}=b^{m}=1,a^{-1}ba=b^{-1}\rangle$ for $m\geq 3$, $n\geq 2$ and the group $G(p,m,n)$ presented by $\langle a,b:a^{p^{m}}=b^{p^{m}}=[a,b]^{p}=1,[a,[a,b]]=[b,[a,b]]=1\rangle$ for any prime $p$ and $m,n\geq 1$. In Section 2.2, we consider commuting conjugacy class graphs of the groups $G$ when $G/Z(G)$ is isomorphic to $D_{2m}$, $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, Frobenious group of order $pq$ and $p^{2}q$ or any group of order $p^{3}$ for any two primes $p$ and $q$.
@@
The commuting conjugacy class graph (in short CCC-graph) of $G$ is a simple undirected graph whose vertex set is the set of conjugacy classes of non-central elements of $G$ and two vertices $x^{G}$ (conjugacy class containing $x\in G$) and $y^{G}$ are adjacent if and only if there exists $x^{\prime}\in x^{G}$ and $y^{\prime}\in y^{G}$ such that $x^{\prime}y^{\prime}=y^{\prime}x^{\prime}$. We write $\mathcal{CCC}(G)$ to denote the CCC-graph of $G$. This graph was introduced by Herzog et al. [17] in 2009. In 2016, Mohammadian et al. [13] have characterized finite groups $G$ such that $\mathcal{CCC}(G)$ is triangle-free. In [26, 27], Salahshour and Ashrafi obtained structures of $\mathcal{CCC}(G)$ of various families of finite CA-groups. Salahshour also obtained $\mathcal{CCC}(G)$ for the groups whose central quotient is isomorphic to a dihedral group [25]. Recently, Rezaei et al. [23] obtained the structures of $\mathcal{CCC}(G)$ when the central quotient of $G$ is isomorphic to a Frobenious group of order $pq$ or $p^{2}q$ for any two primes $p$ and $q$. Various spectral aspects of $\mathcal{CCC}(G)$ have been considered in [4, 6, 11, 12]. In [5], genus of $\mathcal{CCC}(G)$ have been computed and determined several groups such that their CCC-graphs are planar, toroidal, double-toroidal or triple-toroidal. A survey on CCC-graphs of groups can be found in [7].
@@ 2. Zagreb indices of CCC-graphs
Let $\Gamma$ be the disjoint union of the graphs $\Gamma_{1},\Gamma_{2},\dots,\Gamma_{n}$. If $\Gamma_{i}=l_{i}K_{m_{i}}$ for $i=1,2,\dots,k$, where $K_{m_{i}}$’s are complete graphs on $m_{i}$ vertices and $l_{i}K_{m_{i}}$ is the disjoint union of $l_{i}$ copies of $K_{m_{i}}$, then
@@ 2.1. Zagreb indices of CCC-graphs of certain well-known finite groups
In this section, we consider the CCC-graphs for the groups $D_{2m},Q_{4m},V_{8m},SD_{8m},U_{(n,m)}$ and $G(p,m,n)$ and compute their Zagreb indices to finally check if they satisfy Hansen-Vukičević conjecture. The structures of the CCC-graphs of the above mentioned groups are obtained in [27]. However, while obtaining the structure of CCC-graph of $U_{(n,m)}$ in [27] one case was missed which has been pointed out by Jannat and Nath in [10].
@@
We have $\mathcal{CCC}(G)=K_{\frac{m-1}{2}}\sqcup K_{1}$. As such, $|v(\mathcal{CCC}(G))|=\frac{m-1}{2}+1=\frac{m+1}{2}$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{m-1}{2}}{2}=\frac{(m-1)(m-3)}{8}$. Therefore, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\frac{m-1}{2}\left(\frac{m-1}{2}-1\right)^{2}+0=\frac{(m-1)(m-3)^{2}}{8}$
and
$M_{2}(\mathcal{CCC}(G))$ $=\frac{\frac{m-1}{2}\left(\frac{m-1}{2}-1\right)^{3}}{2}+0=\frac{(m-1)(m-3)^{3}}{32}.$
Now
@@
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{(m-3)^{2}}{4}\left(1-\frac{m-1}{m+1}\right)$
$=\frac{(m-3)^{2}}{4}\left(\frac{2}{m+1}\right)\geq 0$
@@
We have $\mathcal{CCC}(G)=K_{\frac{m}{2}-1}\sqcup 2K_{1}$. As such, $|v(\mathcal{CCC}(G))|=\frac{m}{2}-1+2=\frac{m+2}{2}$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{m}{2}-1}{2}=\frac{(m-2)(m-4)}{8}$. Therefore, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\left(\frac{m}{2}-1\right)\left(\frac{m}{2}-1-1\right)^{2}+0=\frac{(m-2)(m-4)^{2}}{8}$
and
$M_{2}(\mathcal{CCC}(G))$ $=\frac{\left(\frac{m}{2}-1\right)\left(\frac{m}{2}-1-1\right)^{3}}{2}+0=\frac{(m-2)(m-4)^{3}}{32}.$
Now
@@
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{(m-4)^{2}}{4}\left(1-\frac{m-2}{m+2}\right)$
$=\frac{(m-4)^{2}}{4}\left(\frac{4}{m+2}\right)\geq 0$
@@
We have $\mathcal{CCC}(G)=K_{\frac{m}{2}-1}\sqcup K_{2}$. As such, $|v(\mathcal{CCC}(G))|=\frac{m}{2}-1+2=\frac{m+2}{2}$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{m}{2}-1}{2}+1=\frac{(m-2)(m-4)}{8}+1$. Therefore, by Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))$ $=\left(\frac{m}{2}-1\right)\left(\frac{m}{2}-1-1\right)^{2}+2(2-1)^{2}=\frac{(m-2)(m-4)^{2}}{8}+2$
and
$M_{2}(\mathcal{CCC}(G))$ $=\frac{\left(\frac{m}{2}-1\right)\left(\frac{m}{2}-1-1\right)^{3}}{2}+\frac{2(2-1)^{3}}{2}=\frac{(m-2)(m-4)^{3}}{32}+1.$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}$ $-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{\frac{(m-2)(m-4)^{3}}{32}+1}{\frac{(m-2)(m-4)}{8}+1}-\frac{\frac{(m-2)(m-4)^{2}}{8}+2}{\frac{m+2}{2}}$
$=\frac{1}{4}\left(\frac{4(m-2)(m-4)(m^{2}-10m+20)+32(m+2)-128}{(m+2)(m-2)(m-4)+8(m+2)}\right)$
$=\frac{m^{4}-16m^{3}+88m^{2}-200m+160+8(m-2)}{(m+2)(m-2)(m-4)+8(m+2)}$
@@
We have $\mathcal{CCC}(G)=K_{m-1}\sqcup 2K_{1}$. As such, $|v(\mathcal{CCC}(G))|=m-1+2=m+1$ and $|e(\mathcal{CCC}(G))|=\binom{m-1}{2}=\frac{(m-1)(m-2)}{2}$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))=(m-1)(m-1-1)^{2}+0=(m-1)(m-2)^{2}$
and
$M_{2}(\mathcal{CCC}(G))=\frac{(m-1)(m-1-1)^{3}}{2}+0=\frac{(m-1)(m-2)^{3}}{2}.$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{\frac{(m-1)(m-2)^{3}}{2}}{\frac{(m-1)(m-2)}{2}}-\frac{(m-1)(m-2)^{2}}{m+1}$
$=(m-2)^{2}-\frac{(m-1)(m-2)^{2}}{m+1}$
$=\frac{(m-2)^{2}(m+1-m+1)}{m+1}=\frac{2(m-2)^{2}}{m+1}\geq 0$
@@
We have $\mathcal{CCC}(G)=K_{m-1}\sqcup K_{2}$. As such, $|v(\mathcal{CCC}(G))|=m-1+2=m+1$ and $|e(\mathcal{CCC}(G))|=\binom{m-1}{2}+1=\frac{(m-1)(m-2)}{2}+1$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))=(m-1)(m-1-1)^{2}+2(2-1)^{2}=(m-1)(m-2)^{2}+2$
and
$M_{2}(\mathcal{CCC}(G))=\frac{(m-1)(m-1-1)^{3}}{2}+\frac{2(2-1)^{3}}{2}=\frac{(m-1)(m-2)^{3}}{2}+1.$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{\frac{(m-1)(m-2)^{3}}{2}+1}{\frac{(m-1)(m-2)}{2}+1}-\frac{(m-1)(m-2)^{2}+2}{m+1}$
$=\frac{(m-1)(m-2)^{3}+2}{(m-1)(m-2)+2}-\frac{(m-1)(m-2)^{2}+2}{m+1}$
$=\frac{m^{4}-8m^{3}+22m^{2}-24m+9}{(m+1)(m-1)(m-2)+2(m+1)}$
$=\frac{m^{2}(m^{2}-8m+14)+8m(m-3)+9}{(m+1)(m-1)(m-2)+2(m+1)}:=\frac{f(m)}{g(m)}.$
@@
We have $\mathcal{CCC}(G)=K_{2m-1}\sqcup 2K_{1}$. As such, $|v(\mathcal{CCC}(G))|=2m-1+2=2m+1$ and $|e(\mathcal{CCC}(G))|=\binom{2m-1}{2}+0=\frac{(2m-1)(2m-2)}{2}$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))=(2m-1)(2m-1-1)^{2}+2(1-1)^{2}=(2m-1)(2m-2)^{2}$
and
$M_{2}(\mathcal{CCC}(G))=\frac{(2m-1)(2m-1-1)^{3}}{2}+\frac{2(1-1)^{3}}{2}=\frac{(2m-1)(2m-2)^{3}}{2}.$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{\frac{(2m-1)(2m-2)^{3}}{2}}{\frac{(2m-1)(2m-2)}{2}}-\frac{(2m-1)(2m-2)^{2}}{2m+1}$
$=(2m-2)^{2}-\frac{(2m-1)(2m-2)^{2}}{2m+1}$
$=\frac{(2m-2)^{2}(2m+1-2m+1)}{2m+1}=\frac{2(2m-2)^{2}}{2m+1}\geq 0$
@@
We have $\mathcal{CCC}(G)=K_{2m-2}\sqcup K_{4}$. As such, $|v(\mathcal{CCC}(G))|=2m-2+4=2m+2$ and $|e(\mathcal{CCC}(G))|=\binom{2m-2}{2}+\binom{4}{2}=\frac{(2m-2)(2m-3)}{2}+6=(m-1)(2m-3)+6$. By Theorem 2.1, we have
$M_{1}(\mathcal{CCC}(G))=(2m-2)(2m-2-1)^{2}+4(4-1)^{2}=(2m-2)(2m-3)^{2}+36$
and
$M_{2}(\mathcal{CCC}(G))=\frac{(2m-2)(2m-2-1)^{3}}{2}+\frac{4(4-1)^{3}}{2}=(m-1)(2m-3)^{3}+54.$
Therefore
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{(m-1)(2m-3)^{3}+54}{(m-1)(2m-3)+6}-\frac{(2m-2)(2m-3)^{2}+36}{2m+2}$
$=\frac{(m-1)(2m-3)^{3}+54}{(m-1)(2m-3)+6}-\frac{(m-1)(2m-3)^{2}+18}{m+1}$
$=\frac{8m^{4}-56m^{3}+120m^{2}-72m}{(m+1)(m-1)(2m-3)+6(m+1)}$
$=\frac{8m^{3}(m-7)+4m(30m-18)}{(m+1)(m-1)(2m-3)+6(m+1)}:=\frac{f(m)}{g(m)}.$

### v2

# Commuting Conjugacy Class Graphs of Finite Groups and the Hansen-Vukičevi’c Conjecture

@@ Abstract.
In this work, we compute the first and second Zagreb indices for the commuting conjugacy class graphs associated with finite groups. We identify multiple classes of finite groups whose commuting conjugacy class graphs are shown to satisfy the Hansen-Vukičević conjecture. Specifically, we prove that the conjecture holds for the commuting conjugacy class graphs of dihedral groups ($D_{2m}$), dicyclic groups, semidihedral groups, and various other two-generator groups. Moreover, we examine the case where the quotient $G/Z(G)$ is isomorphic to $D_{2m}$, $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, a Frobenius group of order $pq$ or $p^{2}q$, or any group of order $p^{3}$, for primes $p$ and $q$. In each of these cases, we demonstrate that the corresponding commuting conjugacy class graph satisfies the Hansen-Vukičević conjecture.
@@
*Correspondence: rajatkantinath@yahoo.com (R. K. Nath); yilun.shang@northumbria.ac.uk (Y. Shang)
@@ 1. Introduction
Let $\mathcal{G}$ denote the collection of all graphs. A topological index for a graph $\Gamma\in\mathcal{G}$ is expressed as $T(\Gamma)$, where $T$ is a real-valued function mapping from $\mathcal{G}$ to $\mathbb{R}$. Originally, such indices were introduced to characterize various chemical attributes of molecular structures. Since 1947, numerous topological indices have been formulated based on different graph parameters. The earliest of these is the Wiener index, a distance-based measure introduced by Wiener [33]. Within the category of degree-based indices, the Zagreb indices are historically the first and were proposed by Gutman and Trinajstić in 1972 [10]. These indices were initially employed to investigate how the total $\pi$-electron energy of a molecule correlates with its structural features. As mentioned in [23], the Zagreb indices have also been applied in the analysis of molecular complexity, chirality, ZE-isomerism, and heterosystems, among others. Over time, mathematicians have explored the general mathematical characteristics of a variety of topological indices. A comprehensive overview of the mathematical aspects of Zagreb indices is available in the survey article [11]. More recent developments on Zagreb indices can be found in e.g. [14, 24, 25].
Let $\Gamma$ be a simple, undirected graph characterized by the vertex set $v(\Gamma)$ and edge set $e(\Gamma)$. Two well-known degree-based topological indices for such graphs are the first and second Zagreb indices, denoted respectively by $M_{1}(\Gamma)$ and $M_{2}(\Gamma)$. These are formally defined as follows:
$M_{1}(\Gamma)=\sum\limits_{v\in v(\Gamma)}\deg(v)^{2}\quad\text{and}\quad M_{2}(\Gamma)=\sum\limits_{uv\in e(\Gamma)}\deg(u)\deg(v),$
where $\deg(v)$ denotes the degree of a vertex $v$, that is, the number of edges incident to it. In an effort to explore the relationship between these two indices, Hansen and Vukičević [12] proposed a conjecture in 2007, which compares the behavior of $M_{1}(\Gamma)$ and $M_{2}(\Gamma)$ for various classes of graphs.
@@
(Hansen-Vukičević Conjecture) Assume that $\Gamma$ is a simple finite graph. We have the following inequality.
@@
It was demonstrated in [12] that the conjecture does not hold in general, as evidenced by the counterexample $\Gamma=K_{1,5}\sqcup K_{3}$. Nevertheless, Hansen and Vukičević [12] verified that Conjecture 1.1 remains valid within the class of chemical graphs. The validity of the conjecture has also been established for trees in [31], with equality in equation (1) occurring specifically when $\Gamma$ is a star graph. For connected unicyclic graphs, the conjecture was confirmed in [18], where equality is attained in the case of cycle graphs. A more detailed examination of equality conditions in (1) was conducted in [32], and a broader survey on the comparison of Zagreb indices is available in [19].
Despite these advances, the complete classification of graphs that either satisfy or violate Conjecture 1.1 remains an open area of investigation. Recently, Das et al. [8] identified various finite non-abelian groups whose commuting or non-commuting graphs fulfill the Hansen-Vukičević conjecture. Similarly, Das and Nath [9] provided several examples of finite groups whose super commuting graphs also conform to the conjecture.
In this paper, we focus on the commuting conjugacy class graphs of finite groups. Specifically, we compute their first and second Zagreb indices and identify several families of finite groups for which these graphs satisfy the Hansen-Vukičević conjecture. Section 2.1 is devoted to analyzing the commuting conjugacy class graphs of certain group families. These include the dihedral groups $D_{2m}$ defined by the presentation $\langle a,b:a^{m}=b^{2}=1,bab^{-1}=a^{-1}\rangle$ for $m\geq 3$; the dicyclic groups $Q_{4m}$ presented as $\langle a,b:a^{2m}=1,b^{2}=a^{m},bab^{-1}=a^{-1}\rangle$ for $m\geq 2$; and the semidihedral groups $SD_{8m}$ with presentation $\langle a,b:a^{4m}=b^{2}=1,bab=a^{2m-1}\rangle$ for $m\geq 2$.
In addition to these, we examine the group $V_{8m}$ given by $\langle a,b:a^{2m}=b^{4}=1,ba=b^{-1}a^{-1},b^{-1}a=a^{-1}b\rangle$ for $m\geq 1$; the group $U_{(n,m)}$ defined by $\langle a,b:a^{2n}=b^{m}=1,a^{-1}ba=b^{-1}\rangle$ for $m\geq 3$, $n\geq 2$; and the group $G(p,m,n)$ with the presentation $\langle a,b:a^{p^{m}}=b^{p^{m}}=[a,b]^{p}=1,[a,[a,b]]=[b,[a,b]]=1\rangle$ for any prime $p$ and integers $m,n\geq 1$. Section 2.2 explores commuting conjugacy class graphs of groups $G$ for which the quotient group $G/Z(G)$ is isomorphic to one of several specific types. These include $D_{2m}$, $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, Frobenius groups of order $pq$ and $p^{2}q$, and all groups of order $p^{3}$, where $p$ and $q$ are prime numbers.
@@
The commuting conjugacy class graph of a group $G$, abbreviated as the CCC-graph, is defined as a simple undirected graph in which each vertex corresponds to a conjugacy class of a non-central element of $G$. Two distinct vertices, denoted $x^{G}$ and $y^{G}$—where $x^{G}$ represents the conjugacy class containing an element $x\in G$—are joined by an edge if and only if there exist representatives $x^{\prime}\in x^{G}$ and $y^{\prime}\in y^{G}$ such that $x^{\prime}y^{\prime}=y^{\prime}x^{\prime}$. This graph is denoted by $\mathcal{CCC}(G)$ and was first introduced by Herzog et al. [13] in 2009. Subsequent research has explored various structural properties of $\mathcal{CCC}(G)$. In 2016, Mohammadian et al. [21] classified all finite groups $G$ for which the CCC-graph $\mathcal{CCC}(G)$ is triangle-free. Further developments were made in [29, 30], where Salahshour and Ashrafi studied the structures of $\mathcal{CCC}(G)$ for several families of finite CA-groups. Additional work by Salahshour in [28] examined CCC-graphs in cases where the quotient group $G/Z(G)$ is isomorphic to a dihedral group. More recently, Rezaei et al. [26] determined the structure of $\mathcal{CCC}(G)$ when $G/Z(G)$ corresponds to a Frobenius group of order $pq$ or $p^{2}q$, where $p$ and $q$ are distinct prime numbers. Spectral characteristics of CCC-graphs have also been explored in a series of papers, including [4, 6, 16, 17]. In [5], the genus of $\mathcal{CCC}(G)$ was computed, and various families of groups were identified whose CCC-graphs are planar, toroidal, double-toroidal, or even triple-toroidal. A comprehensive overview of the CCC-graph and its applications to group theory is available in the survey article [7].
@@ 2. Zagreb indices of CCC-graphs
Suppose that $\Gamma$ is the disjoint union of the graphs $\Gamma_{1},\Gamma_{2},\dots,\Gamma_{n}$. Let $\Gamma_{i}=l_{i}K_{m_{i}}$ for $i=1,2,\dots,k$, where $K_{m_{i}}$’s are complete graphs on $m_{i}$ vertices and $l_{i}K_{m_{i}}$ is the disjoint union of $l_{i}$ copies of $K_{m_{i}}$. We obtain
@@ 2.1. Zagreb indices of CCC-graphs of certain well-known finite groups
In this section, we consider the CCC-graphs for the groups $D_{2m},Q_{4m},V_{8m},SD_{8m},U_{(n,m)}$ and $G(p,m,n)$ and compute their Zagreb indices to finally check if they satisfy Hansen-Vukičević conjecture. The structures of the CCC-graphs of the above mentioned groups are obtained in [30]. However, while obtaining the structure of the CCC-graph of $U_{(n,m)}$ in [30], one case was missed, which has been pointed out by Jannat and Nath in [15].
@@
We have $\mathcal{CCC}(G)=K_{\frac{m-1}{2}}\sqcup K_{1}$. Therefore, by Theorem 2.1, we get the expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$.
Note that $|v(\mathcal{CCC}(G))|=\frac{m-1}{2}+1=\frac{m+1}{2}$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{m-1}{2}}{2}=\frac{(m-1)(m-3)}{8}$. Therefore,
@@
Hence,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{(m-3)^{2}}{4}\left(\frac{2}{m+1}\right)\geq 0$
@@
We have $\mathcal{CCC}(G)=K_{\frac{m}{2}-1}\sqcup 2K_{1}$. Therefore, by Theorem 2.1, we get the expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$.
Note that $|v(\mathcal{CCC}(G))|=\frac{m}{2}-1+2=\frac{m+2}{2}$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{m}{2}-1}{2}=\frac{(m-2)(m-4)}{8}$. Therefore,
@@
Hence,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{(m-4)^{2}}{4}\left(\frac{4}{m+2}\right)\geq 0$
@@
We have $\mathcal{CCC}(G)=K_{\frac{m}{2}-1}\sqcup K_{2}$. Therefore, by Theorem 2.1, we get the expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$.
Note that $|v(\mathcal{CCC}(G))|=\frac{m}{2}-1+2=\frac{m+2}{2}$ and $|e(\mathcal{CCC}(G))|=\binom{\frac{m}{2}-1}{2}+1=\frac{(m-2)(m-4)}{8}+1$. Therefore,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=\frac{(m-2)(m-4)^{3}+32}{4((m-2)(m-4)+8)}\text{ and }\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{(m-2)(m-4)^{2}+16}{4(m+2)}.$
Hence,
@@
We have $\mathcal{CCC}(G)=K_{m-1}\sqcup 2K_{1}$. Therefore, by Theorem 2.1, we get the expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$.
Note that $|v(\mathcal{CCC}(G))|=m-1+2=m+1$ and $|e(\mathcal{CCC}(G))|=\binom{m-1}{2}=\frac{(m-1)(m-2)}{2}$. Therefore,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=(m-2)^{2}\text{ and }\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{(m-1)(m-2)^{2}}{m+1}.$
Hence,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{2(m-2)^{2}}{m+1}\geq 0$
@@
We have $\mathcal{CCC}(G)=K_{m-1}\sqcup K_{2}$. Therefore, by Theorem 2.1, we get the expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$.
Note that $|v(\mathcal{CCC}(G))|=m-1+2=m+1$ and $|e(\mathcal{CCC}(G))|=\binom{m-1}{2}+1=\frac{(m-1)(m-2)}{2}+1$. Therefore,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=\frac{(m-1)(m-2)^{3}+2}{(m-1)(m-2)+2}\text{ and }\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{(m-1)(m-2)^{2}+2}{m+1}.$
Hence,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{m^{2}(m^{2}-8m+14)+8m(m-3)+9}{(m+1)(m-1)(m-2)+2(m+1)}:=\frac{f(m)}{g(m)}.$
@@
We have $\mathcal{CCC}(G)=K_{2m-1}\sqcup 2K_{1}$. Therefore, by Theorem 2.1, we get the expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$.
Note that $|v(\mathcal{CCC}(G))|=2m-1+2=2m+1$ and $|e(\mathcal{CCC}(G))|=\binom{2m-1}{2}+0=\frac{(2m-1)(2m-2)}{2}$. Therefore,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=(2m-2)^{2}\text{ and }\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{(2m-1)(2m-2)^{2}}{2m+1}.$
Hence,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{2(2m-2)^{2}}{2m+1}\geq 0$
@@
We have $\mathcal{CCC}(G)=K_{2m-2}\sqcup K_{4}$. Therefore, by Theorem 2.1, we get the expressions for $M_{1}(\mathcal{CCC}(G))$ and $M_{2}(\mathcal{CCC}(G))$.
Note that $|v(\mathcal{CCC}(G))|=2m-2+4=2m+2$ and $|e(\mathcal{CCC}(G))|=\binom{2m-2}{2}+\binom{4}{2}=\frac{(2m-2)(2m-3)}{2}+6=(m-1)(2m-3)+6$. Therefore,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}=\frac{(m-1)(2m-3)^{3}+54}{(m-1)(2m-3)+6}\text{ and }\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}=\frac{(m-1)(2m-3)^{2}+18}{m+1}.$
Hence,
$\frac{M_{2}(\mathcal{CCC}(G))}{|e(\mathcal{CCC}(G))|}-\frac{M_{1}(\mathcal{CCC}(G))}{|v(\mathcal{CCC}(G))|}$ $=\frac{8m^{3}(m-7)+4m(30m-18)}{(m+1)(m-1)(2m-3)+6(m+1)}:=\frac{f(m)}{g(m)}.$